OpenStudy (anonymous):

A roll of wallpaper is 10 meters long and 0.5 meters wide. Assuming that there is no wastage, how many rolls of wallpaper would Peter require to cover the two walls?

OpenStudy (anonymous):

two balls of what dimension?

OpenStudy (anonymous):

Let me get a picture

OpenStudy (anonymous):

thats a cube not a ball -.-

OpenStudy (anonymous):

The question is asking how many "rolls or balls as you say" will cover the 2 walls if the "rolls" are 10 inches long and 0.5 inches wide

OpenStudy (anonymous):

Area of 2 walls = 2 * ( 5 * 8) = 89 in² Area of 1 roll = 10 * .5 = 5 in² =>Number of rolls: = 89/ 5 ~ 18 rolls

OpenStudy (anonymous):

Isn't 5 x8 = 40?